OpenStudy (anonymous):

Solve the inequality and represent your answer in set builder notation. Be sure to show all of your work x/2 + 6 < 10

OpenStudy (anonymous):

Where are you stucked?

OpenStudy (anonymous):

x/2 + 6 < 10 = x/2 < 10 - 6 = x < 2(10 - 6) And then distribute